APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday October 29, 2004 (MJD 53308) ---=== OBSERVING TEAM ===--- Mike Harvanek Support: Jon Brinkmann Craig Loomis ---=== OBSERVING PLAN ===--- Spectro monthly checkout. ---=== OBSERVING SUMMARY ===--- All tape drives including spectro, PT and the spare were cleaned. A short doghouse bias showed the imager and DA to be functioning properly. Spectro monthly checkout was performed. A lot of positioning (both left/right and wavelength) was needed to get things into spec. After the dithered flats, collimator motor C failed to return to the proper position for BOTH sp1 and sp2. This was corrected before proceeding. See the log for more details. The ODH sensor in the enclosure went off multiple (3?) times tonight. The alarms lasted from 15-20 seconds to a couple of minutes. Lots of fiducial too large to correct errors. See the log for details. ---=== OBSERVING LOG ===--- Monthly checkout (night 1): --------------------------- All tape drives including spectro, PT and the spare were cleaned. The spare drive was doing some weird things. Craig Loomis was notified and is aware of the problem. A short imager doghouse bias was taken and the imager and DA checked out okay. Spectro monthly checkout was performed. Details follow. starting sp1 Coll_motor_A 0 Coll_motor_B -18 Coll_motor_C -1 starting sp2 Coll_motor_A 4 Coll_motor_B -3 Coll_motor_C -1 im> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 6.26714 tpm_TM_TelairAvg.VALA 5.45 Primary and secondary within +/- 2 degrees of ambient. Adjusting the left to right positioning: sp1> mechMoveColl 6400 6400 -6400 sp2> mechMoveColl 2300 2300 -2300 Fine tuning: sp1> mechMoveColl 200 200 -200 sp2> mechMoveColl 200 200 -200 and again: sp1> mechMoveColl 200 200 -200 sp2> mechMoveColl 200 200 -200 Moving back to the previous position for sp1: sp1> mechMoveColl -200 -200 200 Adjusting the wavelength positioning: sp1> mechMoveColl -5000 5000 0 sp2> mechMoveColl 12500 -12500 0 Too far; moving halfway back: sp1> mechMoveColl 2500 -2500 0 sp2> mechMoveColl -6300 6300 0 Fine tuning for sp1: sp1> mechMoveColl 700 -700 0 While moving the rotator to 45 degrees, the ODH sensor went off briefly (15-20 seconds). It went off again shortly thereafter for about the same duration. Spectro collimation: rotator moved to 45 degrees im> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 5.05 tpm_TM_TelairAvg.VALA 4.465 Primary and secondary within +/- 2 degrees of ambient. current sp1 Coll_motor_A 4800 Coll_motor_B 8373 Coll_motor_C -6599 current sp2 Coll_motor_A 8900 Coll_motor_B -3502 Coll_motor_C -2701 sp1> mechPiston 3476 no change needed for sp2 final sp1 Coll_motor_A 8276 Coll_motor_B 11859 Coll_motor_C -3125 final sp2 Coll_motor_A 8900 Coll_motor_B -3502 Coll_motor_C -2701 Dithered flats are frames 29034 to 29040. Arcs after dithered flats are frames 29041 and 29042. Note that after the dithered flats and arcs were completed, collimator motor C for sp1 AND sp2 did NOT return to its proper position: sp1 Coll_motor_A 8290 Coll_motor_B 11840 Coll_motor_C 200 --> difference of -3325 sp2 Coll_motor_A 8899 Coll_motor_B -3508 Coll_motor_C -300 --> difference of -2401 The proper positions were restored with sp1> mechMoveColl 0 0 -3325 and sp2> mechMoveColl 0 0 -2401 Another flat and arc (frames 29043 and 29044) and another set of Hartmann exposures (frames 29045 and 29046) were taken after this to make sure that all was well. The ODH sensor went off again for about a minute or two. The 25 biases are frames 29047 to 29071. Frame 29060 was visually inspected (all 4 cameras) and nothing unusual was observed. The histogram plot and the percentiles for this bias have been printed out and put in the spectro monthly checkout binder. The percentiles for this frame are also shown below: Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00029060.fit -5.5 -4.4 -3.1 2.4 7.1 9.2 10.5 sdR-b1-00029060.fit -5.5 -4.4 -3.1 2.4 7.1 9.2 10.5 sdR-b2-00029060.fit -9.1 -6.8 -5.2 1.2 7.8 9.4 11.7 sdR-b2-00029060.fit -9.1 -6.8 -5.2 1.2 7.8 9.4 11.7 sdR-r1-00029060.fit -7.5 -5.2 -3.9 1.7 7.1 8.6 10.9 sdR-r1-00029060.fit -7.5 -5.2 -3.9 1.7 7.1 8.6 10.9 sdR-r2-00029060.fit -7.8 -5.7 -4.6 0.6 5.9 7.6 9.6 sdR-r2-00029060.fit -7.8 -5.7 -4.6 0.6 5.9 7.6 9.6 The 3 darks are frames 29072 to 29074. The histogram plot and the percentiles for dark frame 29073 have been printed out and put in the spectro monthly checkout binder. The percentiles for this frame are also shown below: Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00029073.fit -3.0 -0.9 0.4 7.0 12.8 14.9 17.2 sdR-b1-00029073.fit -3.0 -0.9 0.4 7.0 12.8 14.9 17.2 sdR-b2-00029073.fit -5.1 -2.5 -1.2 5.7 13.3 16.0 18.6 sdR-b2-00029073.fit -5.1 -2.5 -1.2 5.7 13.3 16.0 18.6 sdR-r1-00029073.fit -4.8 -2.5 -1.4 4.5 11.3 12.5 15.5 sdR-r1-00029073.fit -4.8 -2.5 -1.4 4.5 11.3 12.5 15.5 sdR-r2-00029073.fit -5.6 -3.6 -1.6 4.4 10.4 12.4 14.4 sdR-r2-00029073.fit -5.6 -3.6 -1.6 4.4 10.4 12.4 14.4 Lots of fiducials too large to correct tonight. Most of them were in altitude but two or three were in azimuth which is weird because I didn't move in az tonight. I tried fiducializing the alt axis twice but it didn't help. Hopefully this problem will go away once we get the new fiducial tables implemented (tomorrow I hope). That's all folks. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 4919 02:53Z 03:10Z 100 O 151.23 155.46 41 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): YES QA Procedures Done (y/n): N/A UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 29017, plate -9999 ------- 02:13 29017 0.0 bias ----- sequence 29018, plate 1865 ------- 03:14 29018 10.0 flat test - l/r positioning 03:42 29019 10.0 flat test - l/r positioning 03:57 29020 10.0 flat test - l/r positioning 04:08 29021 10.0 flat test - l/r positioning 04:18 29022 2.0 arc test - wavelength positioning 04:37 29023 2.0 arc test - wavelength positioning 04:53 29024 2.0 arc test - wavelength positioning 05:06 29025 2.0 arc test - wavelength positioning 05:18 29026 10.0 flat test - final positioning 05:32 29027 2.1 arc test - collimation 05:34 29028 2.1 arc test - collimation 05:45 29029 10.0 flat bad - forgot to turn on lamps 05:48 29030 10.0 flat good 05:52 29031 2.0 arc good 05:56 29032 2.1 arc test - collimation 05:58 29033 2.1 arc test - collimation 06:14 29034 10.0 flat dithered flat 06:18 29035 10.0 flat dithered flat 06:24 29036 10.0 flat dithered flat 06:27 29037 10.0 flat dithered flat 06:31 29038 10.0 flat dithered flat 06:35 29039 10.0 flat dithered flat 06:39 29040 10.0 flat dithered flat 06:43 29041 2.0 arc arc after dithered flat 06:46 29042 2.0 arc arc after dithered flat 07:01 29043 10.0 flat good 07:03 29044 2.0 arc good 07:14 29045 2.1 arc test - collimation 07:17 29046 2.1 arc test - collimation 07:26 29047 0.0 bias 07:28 29048 0.0 bias 07:31 29049 0.0 bias 07:33 29050 0.0 bias 07:35 29051 0.0 bias 07:37 29052 0.0 bias 07:39 29053 0.0 bias 07:41 29054 0.0 bias 07:43 29055 0.0 bias 07:45 29056 0.0 bias 07:48 29057 0.0 bias 07:50 29058 0.0 bias 07:52 29059 0.0 bias 07:54 29060 0.0 bias 07:56 29061 0.0 bias 07:59 29062 0.0 bias 08:01 29063 0.0 bias 08:03 29064 0.0 bias 08:05 29065 0.0 bias 08:07 29066 0.0 bias 08:09 29067 0.0 bias 08:11 29068 0.0 bias 08:13 29069 0.0 bias 08:15 29070 0.0 bias 08:17 29071 0.0 bias 09:01 29072 0.0 dark 09:20 29073 0.0 dark 09:37 29074 0.0 dark ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: None Stays: None ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:40Z 36 4 13 345 (NNW) 73 - 01:13Z 36 4 6 358 (NNW) 73 - 01:45Z 36 4 11 7 (N) 74 - 02:17Z 39 2 13 339 (NNW) 45 - 02:50Z 39 6 11 327 (NNW) 50 - 03:22Z 39 6 9 335 (NNW) 44 - 03:57Z 39 10 12 312 (NW) 43 - 04:30Z 40 4 6 282 (WNW) 60 - 05:03Z 39 3 8 264 (W) 43 - 05:35Z 39 -2 9 259 (W) 49 - 06:06Z 40 -9 17 259 (W) 34 - 06:40Z 38 -3 13 254 (WSW) 30 - 07:12Z 38 4 9 274 (W) 72 - 07:44Z 38 6 10 274 (W) 65 - 08:17Z 39 7 13 277 (W) 81 - 08:49Z 39 7 14 272 (W) 68 - 09:20Z 39 21 15 278 (W) 230 - 09:54Z 39 25 14 275 (W) 292 - 10:26Z 38 26 14 270 (W) 366 - ---=== TELESCOPE STATUS ===--- 01:45Z Enclosure doors opened, fans and blowers on, louvers opened. 08:45Z Doors closed. 09:45Z Fans and blowers off, louvers closed. Telescope stowed at 30 degrees with cartridge 1 (plate 1865) mounted. Counterweights at 173. Spectro auto-fill connected. Imager dewar 10:35 188 lb, 22 psi Spectro dewar 10:35 270 lb, 24 psi No interlocks bypassed. ---=== SOFTWARE USED ===--- IOP/SOP: v3_142_0 Watcher: v2_29_0 MCP: v5_25_0 TPM: tpm_v2_42_0 AstroDa: v14_47 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v1_95 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 0.5080 0.5250 Axial B 0.5890 0.6020 Axial C 0.8880 0.9050 Trans D -8.6120 -8.6120 Lateral E 10.0203 10.0203 Lateral F 11.5951 11.5951 GALILS Commanded: 4800. -2150. 2600. -6900. -7850. -7900. Actual: 4782. -2150. 2618. -6880. -7862. -7902. SETMIR VALUES PriDesOrient: 0.00 -2.76 1.04 57.00 151.01 PriOrient: 0.00 -3.04 0.88 57.42 149.40 SECONDARY: ---------- Focus: 0.00 Air Temp.: 3.9 Alt.: 29.973517 MIGS TONIGHT NOMINAL Axial A 1.5720 1.5390 Axial B 1.0400 1.0440 Axial C 1.1610 1.1590 Trans D -0.1950 -0.2020 GALILS Commanded: 1619107. 1604789. 1528500. -13800. -9950. Actual: 1619188. 1604973. 1528571. -13820. -9964. SETMIR VALUES SecDesOrient: 1257.00 27.97 -10.00 -90.00 257.89 SecOrient: 1256.91 27.94 -10.03 -89.96 257.52 ---=== PROBLEMS IN DETAIL ===---